Current position analysis

The following data were taken from the balance sheet of Nilo Company at the end of two recent fiscal years:

Current Year Previous Year
Current assets:
  Cash $502,700 $440,000
  Marketable securities 662,100 519,200
  Accounts and notes receivable (net) 815,200 624,800
  Inventories 1,221,800 695,900
  Prepaid expenses 604,200 432,700
  Total current assets $3,806,000 $2,712,600
Current liabilities:
  Accounts and notes payable (short-term) $797,500 $660,000
  Accrued liabilities 302,500 330,000
  Total current liabilities $1,100,000 $990,000
  1. Determine for each year (1) the working capital, (2) the current ratio, and (3) the quick ratio. Round the ratios to one decimal place and the working capital to the nearest dollar.

    Current Year Previous Year
    1. Working capital $ $
    2. Current ratio
    3. Quick ratio
Answer #1
Current Previous
Working capital

= 3,806,000-1,100,000

= 2,706,000

= 2,712,600-990,000

= 1,722,600

Current Ratio

= 3,806,000/1,100,000

= 3.5

= 2,712,600/990,000

= 2.7

Quick Ratio

= 1,980,000/1,100,000

= 1.8

= 1,584,000/990,000

= 1.6